1 – Consider the car-caravan analogy from Section 4 in Chapter 1. In this problem, assume a propagation speed of 70 km/hr and that each toll booth takes 6 seconds to service a car.

a) (7 points) Suppose the caravan of 10 cars begins immediately in front of the first toll booth, travels 50 km to a second toll booth, then another 50 km to a third toll booth, and finally stops immediately after the third tool booth. Thus, they travel a total of 100 km. What is the total end-to-end delay?

b) (3 points) Where is the last car in the caravan after one hour? Your answer must include a distance/specific location, and not only a relative direction
Answer #1

`Hey,

Note: In case of any queries, just comment in box I would be very happy to assist all your queries

a.)

The cars start at front of first toll booth and ends after the third toll booth.

Transmission delay or dtrans at toll booth = 6 seconds at each toll booth for each car.

There are 10 cars. So total transmission delay at each toll booth = 60 seconds = 1 minutes.

Propagation delay or dprop = distance travelled / propagation speed

Total distance = 100 km.

dprop = 100 km/ 70 kmh = (100/70) * 60 minutes = 85.71.

Total end to end delay = dprop + (time taken at each of the three toll booths)

= dprop + 3 * dtrans (because there are 3 booths)

= 85.71 + 1*3 = 88.71 minutes.

b.)

The last car in the caravan leaves the first toll booth after 60 seconds or 1 minutes. Therefore, the distance of the last car from the first toll booth after 60 minutes will be equal to distance travelled by it in (60 - 1) or 59 minutes

59 minutes = (59/60) hours.

Distance travelled = speed * time

= 70 km/hour * (59/60) hours

= 68.83 km.

Hence the last car will be 68.83 kilometers away from the first toll booth.

Kindly revert for any queries

Thanks.